2. per year on a given acre. Each farmer chooses whether to devote all acres to producing rye or corn or to produce rye on some of the land and corn on 18 symbol) to plot Francess PPF Francess PPF
media%2Fd89%2Fd89be5ab-843c-422c-8df9-c0
Answer #1

Dmitri: Total production of Rye = 18 x 12 = 216

Total production of Corn = 6 x 12 = 72

Frances; Total production of Rye = 28 x 12 = 336

Total production of Corn = 7 x 12 = 84

108 Dmitnis PPF 72 Francess PPF 12 0 120 180 240 300 00 420 40 540 600 RYE (Bushels)

Absolute Advantage:

Absolute advantage is the ability of a country to produce a good or service at a lower per unit cost as compared to any other country that produces same good or service.

Frances has an absolute advantage in the production of both goods Rye and Corn.

Comparative Advantage: A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

Dmitri; 6 bushels of corn = 18 bushels of Rye

1 bushels of corn = 18/6 = 3 bushels of Rye

Opportunity cost of 1 bushel of corn is 3 bushels of Rye.

Frances; 7 bushels of corn = 28 bushels of Rye

1 bushel of corn = 28/7 = 4 bushels of Rye

Opportunity cost of 1 bushel of corn is 4 bushels of Rye.

Dmitri has lower opportunity cost.

Dmitri has comparative advantage in the production of Corn.

Frances has comparative advantage in the production of Rye.